2. The driving time between Toronto and Kitchener is found to range evenly between 80 minutes to 120 minutes. What is the probability that the drive will take: a) less than 95 minutes? b) More than 95 minutes? c) Between 90 minutes and 110 minutes?
Answer #1

here this is uniform distribution with parameter a=80 and b=120

a)therefore P(X<95)=(x-a)/(b-a)=(95-80)/(120-80)=15/40=3/8

b)

P(X>95)=1-P(X<95)=1-3/8=5/8

c)P(90<X<110)=(110-90)/(120-80)=20/40=1/2